OpenStudy (anonymous):

–2 • (4) – 3 • (–5) = A. 55 B. 7 C. –23 D. –120

OpenStudy (anonymous):

55

OpenStudy (anonymous):

i got b

OpenStudy (anonymous):

because -2*4=-8 -8-3=-11 -11*-5=55

OpenStudy (anonymous):

if u use a caculater